OpenStudy (anonymous):

The formula r=2√5L can be used to approximate the speed r, in mph of a car that left skid marks of length L, in feet. How far will a car skid at 80 mph?

OpenStudy (lalaly):

r=2√5L 80=2√5L 40=√5L sq. both sides 1600=5L L=320

OpenStudy (anonymous):

this is maybe (without exaggeration) the 8th time i have seen this exact same "skid marks" problem. one line class?

OpenStudy (anonymous):

oh and look, it is the Math idol judge problem next. where are these coming from?

jimthompson5910 (jim_thompson5910):

\[\large r=2\sqrt{5L}\] \[\large 80=2\sqrt{5L}\] \[\large \frac{80}{2}=\sqrt{5L}\] \[\large 40=\sqrt{5L}\] \[\large 40^2=5L\] \[\large 1600=5L\] \[\large \frac{1600}{5}=L\] \[\large 320=L\] \[\large L=320\] So the car will skid 320 ft.